A contractor buys 15 yd^2 of nylon carpet and 26 yd^2 of wool carpet for $1085. A second purchase, at the same prices, includes 10 yd^2 of nylon carpet and 20 yd^2 of wool carpet for $790. Find the cost per square yard of the wool carpet
Answer by stanbon(75887)

A contractor buys 15 yd^2 of nylon carpet and 26 yd^2 of wool carpet for $1085. A second purchase, at the same prices, includes 10 yd^2 of nylon carpet and 20 yd^2 of wool carpet for $790. Find the cost per square yard of the wool carpet.
----------------
15n + 22w = 1085
10n + 20w = 790
--------
Modify::
10*15n + 10*22w = 10*1085
15*10n + 15*20w = 15*790
-----------------------------------
Subtract and solve for "w"::
300w - 220w = 15*790-10*1085
80w = 1000
w = $12.50 (cost per yd^2 of wool carpet)
------------
Cheers,
Stan H.
-----------